A right-angled triangle has a perimeter of 120cm. If the hypoteneuse is 50cm, find the length of the other two sides.

Answer by checkley71(8403)

A^2+b^2=c^2 &
A+B+C=120
A+B+50=120
A+B=120-50
A+B=70
A=70-B
Then:
(70-B)^2+B^2=50^2
4900-140B+B^2+B^2=2500
2B^2-140B+4900-2500=0
2B^2-140B+2400=0
2(B^2-70B+1200)=0
2(B-40)(B-30)=0
B-40=0
B=40 ANSWER.
A=70-40
A=30 ANSWER.